OpenStudy (anonymous):

For the distribution, find (a) the mean, (b) the variance, and (c) the standard deviation to the nearest tenth.

OpenStudy (anonymous):

42,46,50,50,52,,54,56

OpenStudy (anonymous):

Am I doing it right.

OpenStudy (anonymous):

@Luigi0210 @FibonacciChick666

OpenStudy (fibonaccichick666):

yea, that's the mean

OpenStudy (anonymous):

what is the standard deviation though? :/

OpenStudy (anonymous):

@AravindG @cwrw238 @happytales @tHe_FiZiCx99 @SithsAndGiggles

OpenStudy (the_fizicx99):

Use the formula :3

OpenStudy (anonymous):

I really can't figure out the standard deviation or the variance... I get like 27.1 for the standard deviation, which is incorrect.

OpenStudy (the_fizicx99):

\(\ \sf \sigma = \sqrt{\Sigma\dfrac{(x - x̄)^2}{n}}\) X_X that took awhile >.>

OpenStudy (anonymous):

so what would I plug in? or what is x equal to??

OpenStudy (anonymous):

an n= ?

OpenStudy (the_fizicx99):

Σ = take the sum of x = each data value x̄ = mean of all the data values n = number of data values

OpenStudy (anonymous):

so x=350?

OpenStudy (the_fizicx99):

No, don't add them. :o this is going to take sometime :3 First calculate the mean, you said the mean was 50. So I'll trust you :3 1. Subtract the mean from each data point. 2. Square the differences you found in step 1 3. Add the squares from step 2 4. Divide the sum from step 3 by the number of data values 5. Find the square root of the number from step 4

OpenStudy (anonymous):

is the drawing I made correct so far?

OpenStudy (the_fizicx99):

Looks blurry to me ;_;

OpenStudy (anonymous):

this is too hard... I might quit lol

OpenStudy (anonymous):

the deviations without being squared are -8,-4,50,2,4,6 correct?

OpenStudy (the_fizicx99):

Honestly it looks hard but it really isn't :O Take your data values, and subtract them with the mean, or 50 42 - 50 = -8 46- 50 = -4 50- 50 = 0 50- 50 = 0 52- 50 = 2 54- 50 = 4 56- 50 = 6 Now square each of those new values, add them up and divide by how many there are and that's it

OpenStudy (anonymous):

it was the 50-50, I was doing 100-50! smh

OpenStudy (the_fizicx99):

Lel, I copied the data you had above, hopefully those were your values and your mean is correct

OpenStudy (anonymous):

so 64+16+0+0+4+16+36/7

OpenStudy (anonymous):

thank you!

OpenStudy (the_fizicx99):

Yes, 19.42 is the Variance, the standard deviation is the square root of that, which is 4.40

OpenStudy (anonymous):

yep , that's what I got! Thanks again!

OpenStudy (the_fizicx99):

You're welcome :>